OpenStudy (anonymous):

cube rt 6xy^3 (times) cube rt 18x^4y

OpenStudy (watchmath):

\(\sqrt[3]{6xy^3}\sqrt[3]{18x^4y}=\sqrt[3]{4\cdot 27\cdot x^3\cdot y^3\cdot x^2\cdot y}=3xy\sqrt[3]{4x^2y}\)